OpenStudy (anonymous):

Hi, need help with this explanation " the square root of an even power of x is obtained by dividing the exponent by 2."

OpenStudy (anonymous):

for example \[\sqrt{x6}\]

OpenStudy (anonymous):

\[\sqrt[b]{x^a}=x^\frac{ a }{ b }\] Where 'a' and 'b' are real numbers

OpenStudy (anonymous):

example. \[\sqrt[2]{x^6}=x ^{\frac{ 6 }{ 2 }}\]

OpenStudy (anonymous):

oh well thank you Donny this makes more sense than what I was getting from the textbook.